"Set up an equation and solve the following problem. How many liters of pure alcohol must be added to 20 liters of a 30% solution to obtain a 90% solution?"
Can I get help with how to work this problem?
Thank you

Answer by josgarithmetic(39617)

(Amount of alcohol)/(amount of total solution) = (concentration of alcohol)
Let v = unknown liters amount of pure alcohol, 100% to add.
Amount of alcohol = v*100+20*30
Amount of resulting solution = v+20
Concentration of alcohol expected for resulting solution = 90%

Form the proper equation and solve for the unknown variable, v.